Tatum Company has four products in its inventory. Information about the December 31, 2018, inventory is as follows:

Product Total Cost Total Net
Realizable Value
101 $ 120,000 $ 100,000
102 90,000 110,000
103 60,000 50,000
104 30,000 50,000


Required:
1. Determine the carrying value of inventory at December 31, 2018, assuming the lower of cost or net realizable value (LCNRV) rule is applied to individual products.
2. Assuming that inventory write-downs are common for Tatum Company, record any necessary year-end adjusting entry.
Answer #1
Concepts and reason

Inventory: It is the goods of the company which can be in forms of raw material, work in progress and finished goods. The company held these goods for sale in the ordinary course of business it shows the capability of the company to manufactured goods in what quantity it is showed in the asset side of the balance sheet.

Fundamentals

Net realizable value: Net realizable value is the expected selling price of inventory minus costs which are to be incurred to prepare the item for sale.

Purchases: Purchase of goods can either be made in cash or on credit from the suppliers. In the case when the goods are purchased on Credit, the company is liable to make its payment at a future date. To record such a liability until the payment is made by the company to the creditor the company maintains an account known as Trade Payables or Accounts Payable.

Cost of Goods Sold: It represents the cost attributable to the production of goods which are sold by the enterprise. It includes direct material cost, direct labor cost which incurred for production excluding indirect cost.

Adjusting entries: Adjusting entries are the entries that are passed at the year-end but before the finalization of the books of accounts. These entries are passed in order to make necessary adjustment of the transactions that involve accruals and deferrals.

1.

Prepare a table to show inventory valuation, using MS Excel as follows:

Valuation of Inventory
Net Realizable
Inventory
value (NRV)
Product
Cost
101
$120.000
$100,000
$100,000
102
$90,000
$110,000

Hence, the carrying value of inventory as on December 31, 2018, is $270,000.

Working Notes:

Compute the total cost of inventory using the equation as follows:

Totalcost=(Costofproduct101+Costofproduct102+Costofproduct103+Costofproduct104)=$120,000+$90,000+$60,000+$30,000=$300,000\begin{array}{c}\\{\rm{Total cost}} = \left( \begin{array}{l}\\{\rm{Cost of product 101}} + {\rm{Cost of product 102}} + \\\\{\rm{Cost of product 103}} + {\rm{Cost of product 104}}\\\end{array} \right)\\\\ = \$ 120,000 + \$ 90,000 + \$ 60,000 + \$ 30,000\\\\ = \$ 300,000\\\end{array}

Hence, the total cost of inventory is $300,000

Compute the total value of inventory using the equation as follows:

TotalInventory=(Valueofproduct101+Valueofproduct102+Valueofproduct103+Valueofproduct104)=$100,000+$90,000+$50,000+$30,000=$270,000\begin{array}{c}\\{\rm{Total Inventory}} = \left( \begin{array}{l}\\{\rm{Value of product 101}} + {\rm{Value of product 102}} + \\\\{\rm{Value of product 103}} + {\rm{Value of product 104}}\\\end{array} \right)\\\\ = \$ 100,000 + \$ 90,000 + \$ 50,000 + \$ 30,000\\\\ = \$ 270,000\\\end{array}

Hence, the total value of inventory is $270,000.

2.

Pass a journal entry to record write-down of inventory as follows:

Post
ref.
Date
Description
31-Dec Cost of goods sold
Inventory
(To record write down of inventory)
Debit($) Credit(s)
$30,000

Working Note:

Compute the inventory write-down using the equation as follows:

Inventorywritedown=TotalCostCarryingvalue=$300,000$270,000=$30,000\begin{array}{c}\\{\rm{Inventory write down}} = {\rm{Total Cost}} - {\rm{Carrying value}}\\\\ = \$ 300,000 - \$ 270,000\\\\ = \$ 30,000\\\end{array}

Hence, inventory will be written down by $30,000.

Ans: Part 1

The carrying value of inventory as on December 31, 2018, is $270,000.

Part 2

Post
ref.
Date
Description
31-Dec Cost of goods sold
Inventory
(To record write down of inventory)
Debit($) Credit(s)
$30,000